An object is projected from the ground with an initial velocity of 20.25 m/s at 50degrees above the horizontal. Find the following. (Show your complete solution.)                               a. Horizontal and vertical components of its initial velocity              b.. Time to reach the maximum height                          c. Range


1
Expert's answer
2022-05-28T08:23:04-0400

Given:

"v_0=20.25\:\rm m/s"

"\theta=50^\circ"

(a)

"v_{0x}=v_0\cos\theta=20.25*\cos50^\circ=13.0\:\rm m/s"

"v_{0y}=v_0\sin\theta=20.25*\sin50^\circ=15.5\:\rm m/s"

(b)

"t_1=\frac{v_{0y}}{g}=\frac{15.5}{9.8}=1.58\rm\: s"

(c)

"R=v_{0x}*2t_1\\
=13.0*2*1.58=41.2\;\rm m"